OpenStudy (anonymous):

lim x->pie/4 (1-tanx)/(sinx-cosx)

OpenStudy (anonymous):

\[\lim_{x\to\pi}\frac{1-\tan x}{\sin x-\cos x}\] step 1) replace tan by sin and cos

OpenStudy (anonymous):

ok then wat do I do

OpenStudy (anonymous):

then things cancel themselves up :)

OpenStudy (anonymous):

so I have (cosx-sinx/cosx)/sinx-cosx

OpenStudy (anonymous):

use the equation editor below to enter your equation. otherwise, I cannot understand

OpenStudy (anonymous):

so then I just cross cancel ?

OpenStudy (anonymous):

ok one min plz

OpenStudy (anonymous):

\[\left[ cosx-sinx/cosx \right]\div \left[ sinx-cosx \right]\]

OpenStudy (anonymous):

so now I just cross cancel?

OpenStudy (anonymous):

\[\lim_{x\to\pi}\Large \frac{-\cancel{(\sin x-\cos x)}\over\cos x}{\cancel{\sin x-\cos x}}\]

OpenStudy (anonymous):

\[=\lim_{x\to\pi}{-1\over\cos x}\]

OpenStudy (anonymous):

then I just plug in \[\pi/4\] for cosx

OpenStudy (anonymous):

I didn't know I could do that within a limit

OpenStudy (anonymous):

thank u very much :)

OpenStudy (anonymous):

you can divide if and only if it is not zero. by definition, as \(x\to\pi/4\), \(x\ne\pi/4\) and hence, \(\sin x\ne\cos x\implies\sin x-\cos x\ne 0\) hence, they can be cancelled. :o

OpenStudy (anonymous):

oh ok thank u :)

OpenStudy (anonymous):

yw.

OpenStudy (anonymous):

uh one question if u don't mind

OpenStudy (anonymous):

shoot it

OpenStudy (anonymous):

um I don't know how u got the negative bcz \[\left( 1-tanx/ sinx-cosx \right)\] \[\left[ 1-sinx / cosx \right] / (sinx-cosx)\] them cross cancel and I get

OpenStudy (anonymous):

\[\left[ 1/cosx \right]\]

OpenStudy (anonymous):

so I don't know where the negative comes from

OpenStudy (anonymous):

\[1-{\sin x\over\cos x}={\cos x-\sin x\over \cos x}\] look the factors you are cancelling!!

OpenStudy (anonymous):

\[(a-b)=-(b-a)\]

OpenStudy (anonymous):

yes then I cancel that out with \[\left[ 1divsinx-cosx \right]\]

OpenStudy (anonymous):

take you time with the equation editor :) atleast I will know what you are saying!

OpenStudy (anonymous):

sorry its my first time

OpenStudy (anonymous):

that is ok. Welcome to Open Study :D

OpenStudy (anonymous):

\[\frac{ cosx-sinx }{ cosx } \times \frac{ 1 }{ sinx-cosx }\] then I cross cancel so where is the negative comeing from ...thanx for the the welcome

OpenStudy (anonymous):

you cannot cancel as yet. they are not same... one says \((a-b)\) the other says \((b-a)\)

OpenStudy (anonymous):

ok yes I understand that

OpenStudy (anonymous):

then, can you make them look same? yes by taking out the "negative" sign

OpenStudy (anonymous):

\[\frac{-(\sin x-\cos x)}{\cos x}\times{1\over \sin x-\cos x}\]

OpenStudy (anonymous):

if I take out the negative sign from \[\left( cosx-sinx \right)\] then wouldn't it be \[-\left( sinx+cosx \right)\]

OpenStudy (anonymous):

ooopsss wait

OpenStudy (anonymous):

I mean it would be \[-\left( cosx+sinx \right)\]

OpenStudy (anonymous):

\[a-b=-(-a)-(b)=-(-a+b)=-(b-a)\]

OpenStudy (anonymous):

you have to flip all the signs. positive become negative and negatives become positives

OpenStudy (anonymous):

oh my gosh I got it im so sorry I confused my self more than necessary hehe

OpenStudy (anonymous):

:) happens to the best of us. This is an indication that you learnt something.

OpenStudy (anonymous):

yup most definitely, thank u for ur help :)

OpenStudy (anonymous):

you are welcome.